a 25ml sample of dilute h2so4 is titrated with 0.0965m naoh so that both protons of h2so4 are neutralized it takes 17.83ml of base to reach end point what is the molarity of h2so4
1
Expert's answer
2020-10-16T08:58:34-0400

Given,

Volume of H2SO4, (V1) = 25ml

Strength of NaOH = 0.0965 m

For the titration purpose Normal strength of NaOH will used

So,

Normally of the NaOH = (0.0965×no. of OHin 1 molicule of NaOH)

=( 0.0965×1 ) N

= 0.0965 N

Normally of the NaOH, S2 =0.0965 N

Volume of NaOH takes,V2 = 17.83ml


Let, the normal strength of H2SO4 = S1 N

For Acid base neutralization reaction,

V1S1 = V2S2

Or, S1 = (V2S2/V1)

= ( 0.0965 N ×17.83 ml )/25 ml

= 0.0688 N


For H2SO4

Normal strength= ( molarity × no. of proton in it )

Or, Normal strength= ( molarity × 2 )

Or, molarity = (normal strength/2)

Or, molarity =( 0.0688/2 ) M

= 0.0344 M


Hence, molarity of H2SO4 = 0.0344 M